# Review on the decoding algorithms for surface codes

### v4

@@ Abstract
Quantum technologies have the potential to solve computationally hard problems that are intractable via classical means. Unfortunately, the unstable nature of quantum information makes it prone to errors. For this reason, quantum error correction is an invaluable tool to make quantum information reliable and enable the ultimate goal of fault-tolerant quantum computing. Surface codes currently stand as the most promising candidates to build near term error corrected qubits given their two-dimensional architecture, a requirement of only local operations, and high tolerance to quantum noise. Decoding algorithms are an integral component of any error correction scheme, as they are tasked with producing accurate estimates of the errors that affect quantum information, so that it can subsequently be corrected. A critical aspect of decoding algorithms is their speed, since the quantum state will suffer additional errors with the passage of time. This poses a connundrum-like tradeoff, where decoding performance is improved at the expense of complexity and viceversa. In this review, a thorough discussion of state-of-the-art surface code decoding algorithms is provided. This work is oriented to readers which may be introducing themselves within the field or willing to further their understanding of the relevant decoders for the surface code, as well as of the state-of-the-art of the field. The core operation of these methods is described along with existing variants that show promise for improved results. In addition, both the decoding performance, in terms of error correction capability, and decoding complexity, are compared. A review of the existing software tools regarding surface code decoding is also provided.
@@ 1 Introduction
Significant progress has been made in the field of quantum computing since Feynman first introduced the idea of computers that use the laws of quantum mechanics in 1982 [1]. Quantum computers leverage the principles of quantum mechanics to accelerate computational processes that are deemed as intractable by means of classical machines, leading to unprecedented results [2]. This technology presents an unprecedented potential to solve problems that are computationally hard since algorithms that present an exponential speedup in terms of computational complexity have been theoretically proposed [3, 4]. In this sense, these novel processors are expected to revolutionize the modern society by improving applications related to cryptography [5, 6], optimization [7, 8, 9], macromolecule design [10, 11, 12] or basic science [13, 14], among others. Recently, quantum supremacy or advantage has been experimentally proven [15, 16, 17, 18, 19], thus, demonstrating that the potential promised by the theory can indeed be realized in physical machines.
@@
There are many possible paths towards the construction of quantum processors such as qubit gate-based quantum computing [20], qudit-based quantum computing [21], measurement based quantum computing [22, 23], quantum annealers [24] or bosonic quantum computing [25], to name a few. Additionally, many physical implementations of these types of technologies like superconducting circuits [26], ion-traps [27] or photons [28] are currently being investigated. In the present article, we will restrict our discussion for qubit gate-based quantum computation, in which the physical elements are realized by discrete two-level systems.
@@
Nevertheless, quantum noise stands as a big obstacle in the way of the groundbreaking promise that quantum technologies offer theoretically [29, 30]. The qubits which constitute quantum computers are prone to suffer from errors, implying that the computations being performed are inaccurate and, therefore, unuseful. The errors that arise whenever computations are being performed via qubits have many origins [30]. For example, quantum gates, which are the quantum analogue of logic gates of classical computation, may introduce unwanted errors to the quantum states being processed due to the fact an over- or under-rotation is being done due to the lack of precision of the elements used for doing such gates [31]. Many of the errors arise due to similar limits in the equipment being used for preparing , controlling or measuring such quantum mechanical systems, and it may seem that dealing with errors suffered by qubits is an engineering problem [32, 33]. However, these elements do also suffer from errors whose origin is their inescapable interaction with their surrounding environment [30, 34, 35]. These types of errors, which are grouped under the term decoherence, constitute a very problematic error source since they always exist. However, dealing with errors is no new feature in the field of computer science [36, 37, 38]. The study of methods for preventing the effects of noise in classical computers is a rich subject area which follows us to the time being, being present in commodities so important as WiFi and other communication devices. Unfortunately, phenomena particular to the quantum mechanics field such as the no-cloning theorem or the effects of measurement in a quantum state prevent several classical methods to be implemented seamlessly in the quantum realm. Luckily, quantum error correction (QEC) was demonstrated to be possible theoretically when the 9-qubit Shor code was proposed in 1995 [29] and generalized afterwards with the theory of Quantum Stabilizer Codes (QSC) by Gottesman in his PhD thesis [39]. The field of quantum error correction has advanced significantly since its earliest formulation and several families of quantum error correction codes (QECC) such as Topological codes [32, 40, 41, 42, 43, 44], Quantum Low-Density Parity Check (QLDPC) codes [45, 46, 47, 48, 49, 50] or Quantum Turbo Codes (QTC) [51, 52, 53], for example, have been proposed.
@@
The basic idea behind quantum error correction can be understood as having the information of a set of qubits (logical qubits) encoded in a greater number of qubits (physical qubits) via a mapping named encoding [30, 39, 54]. In order to estimate the errors suffered from those qubits, a non-destructive measurement (with this we mean that the actual qubits constituting the code are not directly measured) named syndrome measurement is done so that useful information about such error is retrieved [30]. The obtained syndrome is then used so that an error candidate for returning the code to the previous undamaged state is estimated. This process, referred as syndrome decoding depends on the specific code construction [32, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53] and is a critical task for the QECCs to work correctly. Decoding quantum error correction codes is a subtle task due to a quantum unique effect named code degeneracy [46, 54], which refers to the existence of errors that share the same syndrome, but transform the quantum state in an indistinguishable manner. As a consequence of degeneracy, the optimal decoding of stabilizer codes has been proven to belong to the #\#P-complete complexity class, which is computationally much harder than decoding classical linear codes [54, 55]. This high complexity imposes a trade-off between decoding performance and decoding time, as related to complexity of the decoding algorithm, due to the fact that decoders of quantum error correction codes should be fast enough in order to correct the noisy quantum state before it suffers from more errors or decoheres completely [56, 57]. Therefore, the field of constructing more efficient, in terms of speed and correction performance, decoders for stabilizer codes is a very active and relevant field of quantum error correction.
@@
Surface codes are one of the most promising families of codes for constructing primitive fault-tolerant quantum computers in the near-term [32, 40, 58, 59, 60, 61, 62]. This family of topological codes presents the benefit of being implementable in two-dimensional grids of qubits with local syndrome measurements, or check operators in the surface code terminology, while presenting a high tolerance to quantum noise [59]. Considering that many of the physical platforms being considered for constructing quantum processors such as quantum dots, cold atoms or even the mainstream superconducting qubits have architectural restrictions limiting qubit connectivity; surface codes represent a perfect candidate for implementing QEC using those technologies. Recently, major breakthroughs in the field of quantum error correction have been achieved with the first successful experimental implementations of surface codes over superconducting qubit processors [63, 64]. The result by Google Quantum AI is specially relevant due to the fact that it experimentally proves that QEC improves its performance when the distance of the surface code increases [64].
@@
In this sense, designing decoders for surface codes is a really important task for near-term quantum computers. At the time of writing, there are many potential methods in order to perform the inference of the error from the syndrome data, each of them with their own strong and weak points. Generally, the performance versus decoding complexity trade-off stands for those methods and, therefore, each of them is a potential candidate for being the one selected for experimental quantum error correction depending on the specific needs of the system to be error corrected. Due to the fact that surface code decoding is a relevant and timely topic, the main aim of this tutorial is twofold. The first is to provide a compilation and a comprehensive description of the main decoders for the surface code, while the second is to offer a comparison between those methods in terms of decoding complexity and performance. With such scope in mind, we first provide an introductory section describing the basic notions of stabilizer code theory in Section 2 so that surface codes can be introduced in Section 3. We follow by discussing the noise sources in quantum computers and the way they are modelled in Section 4. These sections serve as preliminary background in order to understand the decoding methods that are discussed in the core of the review, Section 5. We describe the most popular surface code decoding algorithms: the Minimum-Weight Perfect Matching (MWPM) decoder, the Union-Find (UF) decoder, the Belief Propagation (BP) decoder and the Tensor Network (TN) decoder. We present their functioning in a comprehensive manner and later discuss not only their performance under depolarizing and biased noise via simulations but also their computational complexity. In addition, we also discuss other decoding methods proposed through the literature for general surface (topological) codes. Also, we review existing and publicly available software implementations of the discussed decoding methods. We then provide a discussion section, Section 6, where we compare the decoders for rotated planar codes and provide an overview of the challenges in decoding surface codes.
@@ 2 Background
In stark contrast, the building block of quantum computers are the elements referred as qubits. Those quantum mechanical systems, named by Benjamin Schumacher in 1995 [65], are two-level quantum systems that allow coherent superpositions of them. In this sense, a qubit can be described as a vector in a two-dimensional complex Hilbert space, $\mathcal{H}_{2}$ [20]:
@@
This and other properties of quantum mechanics such as entanglement [66] allow for a series of advantages in computation (speedups in algorithm complexity [3, 4]), or in communications (supperadditivity of the quantum channel capacity [67, 68, 69, 70]), among others. Nevertheless, such promises are put in question by the inherent noise present in these quantum mechanical systems. As good as it is to make use of quantum unique properties such as superposition or entanglement, quantum noise does also follow different laws and, thus, it is somewhat different to the noise in classical computers/communications. Classical bit noise can be summarized to flips, a logical operation which would turn the 0 into $1$ and vice versa. We refer to Section 4 for understanding quantum noise in general, but as it will be seen there, for the case of a single qubit the noise can be described by means of Pauli channels. The Pauli channel refers to a stochastic model where a qubit can suffer from bit-flips, phase-flips or bit-and-phase-flips each with some probability. With this consideration in mind, these noise effects are described by the elements of the Pauli group [30, 54], $\mathcal{G}_{1}$, whose generators are the Pauli matrices $\langle X,Y,Z\rangle$. Those operators transform the state of an arbitrary qubit as in eq.2 in the following way:
@@
Note that the the $Y$ operator not only does perform a bit-and-phase-flip operation on the arbitrary quantum state but also changes its overall phase. However, neglecting the global phase has no observable physical consequences and, thus, it is often completely ignored from the point of view of quantum error correction [30, 54, 71].
Dealing with these noise processes is a task of vital importance if complex quantum algorithms are meant to be executed reliably. In this context, there are two main approaches to deal with noisy quantum computations: quantum error mitigation (QEM) and quantum error correction (QEC). The two approaches have shown to be complementary with recent papers proposing schemes such as distance scaled zero-noise extrapolation (DSZNE) [72]. QEM attempts to evaluate accurate expectation values of physical observables of interest by using noisy qubits and quantum circuits [72, 73, 74, 75, 76], while the main objective of QEC is to obtain qubits and computations that are “noiseless”. There are many approaches to quantum error correction, but the general idea behind QECCs is to protect the information of a number of qubits $k$, named logical qubits, within a larger number of qubits $n$, named physical qubits in a way that makes the whole system tolerant to a number of errors. Within these QECCs, many lay within the framework named Quantum Stabilizer Coding [39]. Stabilizer codes allow for a direct mapping of many classical methods into the quantum spectrum, making them very useful [30, 39, 54]. Since the surface code belongs to the family of QSCs, it is pertinent to cover the basics of such QEC constructions.
@@ 2.1 Stabilizer Codes
if $E\in\mathcal{S}\subset\mathcal{G}_{n}$. In this sense, there will be different error operators that share the same error syndrome that affects the encoded quantum state in a similar manner. This phenomenon is usually termed as degeneracy. The concept of error degeneracy has the consequence that the Pauli space that represents all possible error operators is not just partitioned into syndrome cosets, but also into degenerate error cosets [^7] Note that degeneracy is somehow different in the entanglement-assisted paradigm [52, 77] [54]. Specifically, the Pauli group is partitioned in $2^{n-k}$ cosets that share error syndrome, and each of those cosets will be partitioned in to $2^{2k}$ cosets that contain $2^{n-k}$ errors that are degenerate among them [67, 54]. How degenerate a quantum code is depends on the difference between the weight of its stabilizer generators and its distance. If $w(M_{k})<<d$, $\forall k\in{0,\ldots,n-k}$, where $M_{k}\in\mathcal{S}$ denotes a stabilizer generator and $w$ denotes the weight, then each logical coset (equivalence class) will contain many operators of the same weight and the code will be highly degenerate. In cases where $w(M_{k})=d$, the code will be non-degenerate.
In summary, the checks give us a partial information of the error operator that corrupted the encoded information. Since the aim of quantum error correction is to recover the noiseless quantum state, an estimate of the channel error must be obtained so that the noisy state can be corrected. The process of estimating the quantum error from the measured syndrome is named decoding. Once a guess of the error, $\hat{E}\in\mathcal{G}_{n}$, is obtained by the decoder, its complex conjugate will be applied to the encoded quantum state. If the estimation turns out to be correct, the noisy quantum state will be successfully corrected since the elements of the Pauli group are unitary. Moreover, if the estimated error is not the exact element of the Pauli group but it belongs to the same degenerate coset, the correction will also be succesful [54]. Finally, whenever the estimated error does not fulfill any of those two cases, the correction operation will result in a non-trivial action on the logical qubits encoded in the state, implying that the error correction method has failed.

### v5

@@ Abstract
Quantum technologies have the potential to solve certain computationally hard problems with polynomial or super-polynomial speedups when compared to classical methods. Unfortunately, the unstable nature of quantum information makes it prone to errors. For this reason, quantum error correction is an invaluable tool to make quantum information reliable and enable the ultimate goal of fault-tolerant quantum computing. Surface codes currently stand as the most promising candidates to build near term error corrected qubits given their two-dimensional architecture, the requirement of only local operations, and high tolerance to quantum noise. Decoding algorithms are an integral component of any error correction scheme, as they are tasked with producing accurate estimates of the errors that affect quantum information, so that they can subsequently be corrected. A critical aspect of decoding algorithms is their speed, since the quantum state will suffer additional errors with the passage of time. This poses a connundrum, where decoding performance is improved at the expense of complexity and viceversa. In this review, a thorough discussion of state-of-the-art decoding algorithms for surface codes is provided. The target audience of this work are both readers with an introductory understanding of the field as well as those seeking to further their knowledge of the decoding paradigm of surface codes. We describe the core principles of these decoding methods as well as existing variants that show promise for improved results. In addition, both the decoding performance, in terms of error correction capability, and decoding complexity, are compared. A review of the existing software tools regarding surface codes decoding is also provided.
@@ 1 Introduction
Significant progress has been made in the field of quantum computing since Feynman first introduced the idea of computers that use the laws of quantum mechanics in 1982 [1]. Quantum computers leverage the principles of quantum mechanics to accelerate computational processes known to be not efficiently solvable by means of classical machines [^1] Efficient in this context means that the problem can be solved in a reasonable amount of time.[2]. Quantum algorithms that offer exponential speedups in terms of computational complexity have been proposed [3, 4], implying that his technology has the unprecedented potential to solve computationally hard problems. In this sense, quantum processors are expected to revolutionize modern society by boosting fields like cryptography [5, 6], optimization [7], macromolecule design [8] or basic science [9], among others. Recent results have demonstrated experimental quantum supremacy or advantage [10, 11, 12, 13, 14], thus, proving that the potential promised by the theoretical quantum algorithms can be realized in physical machines.
@@
There are many possible paths towards the construction of quantum processors such as qubit gate-based quantum computing [15], qudit-based quantum computing [16], measurement based quantum computing [17, 18], quantum annealers [19] or bosonic quantum computing [20], to name a few. Many physical implementations of these types of technologies like superconducting circuits [21], ion-traps [22] or photons [23] are currently being investigated. In the present article, we will restrict our discussion for qubit gate-based quantum computation, in which the physical elements of a quantum computer are realized by discrete two-level systems.
@@
Despite recent advances, quantum noise still stands as a big obstacle to practical quantum computing [24]. The qubits that make up current quantum computers are prone to suffer from errors, leading to inaccurate computations and unreliable results. There are numerous mechanisms by which qubits suffer from errors during computation. For example, quantum gates, which are the quantum analogue of classical logic gates, can cause unwanted errors to the qubits they are applied to due to imprecise implementation (this can be understood as an over- or under-rotation of the qubits targeted by the gate) [25]. Many of the errors arise due to similar imperfections in the equipment being used for preparing, controlling or measuring such quantum mechanical systems, and it may seem that dealing with errors suffered by qubits is an engineering problem [25, 26]. However, qubits also suffer from errors whose origin is their inescapable interaction with their surrounding environment [27, 28, 29]. These types of errors, which are grouped under the term decoherence, represent an unavoidable and hence, problematic error source. However, dealing with errors is not a new requirement in the field of computer science. Methods to prevent and mitigate the effects of noise in classical systems have been a thoroughly studied subject in the past century, and they have enabled commodities as important as WiFi. Unfortunately, phenomena particular to quantum mechanics, such as the no-cloning theorem or the effects of measurement on a quantum state, prevent the seamless implementation of classical methods in the quantum realm. Quantum error correction (QEC) was shown to be feasable (at least theoretically) when the 9-qubit Shor code was proposed in 1995 [24]. The field was soon generalized following this demonstration with the theory of Quantum Stabilizer Codes, introduced by Gottesman in his PhD thesis [30]. The field of quantum error correction has advanced significantly since then and several families of quantum error correction codes (QECC) such as Topological codes [25, 31, 32, 33, 34, 35], Quantum Low-Density Parity Check (QLDPC) codes [36, 37, 38, 39] or Quantum Turbo Codes (QTC) [40] have been proposed.
@@
The core principle of quantum error correction is embedding the information contained in a number of qubits (logical qubits) in a larger number of qubits (physical qubits) via a mapping known as encoding [27, 30, 41]. To estimate the errors suffered by the logical qubits, non-destructive measurements (with this we mean that the actual qubits constituting the code are not directly measured) known as syndrome measurements, are performed [27]. These measurements yield data known as syndromes, which are used to produce an estimate of the error that took place, which, in turn, is used to return the quantum state to the previous undamaged state. This process, referred as syndrome decoding, depends on the specific code construction [25, 32, 33, 34, 35, 36, 37, 38, 39, 40] and is critical for the QECCs to work. Decoding quantum error correction codes is further nuanced by a quantum unique effect known as code degeneracy [37, 41], which refers to the existence of different errors that share the same syndrome (these errors cannot effectively be distinguished by measurements) but transform the quantum state in an indistinguishable manner. As a consequence of degeneracy, the optimal decoding of stabilizer codes falls within the #\#P-complete complexity class, which is computationally harder than decoding classical linear codes [41, 42]. The complexity of the problem imposes a trade-off between decoding performance and decoding time. While decoders for quantum error correction codes need to be as accurate as possible, they should also be fast enough to correct the noisy quantum states before they suffer from more errors or decohere completely [43, 44]. Thus, the design and implementation of efficient decoders for stabilizer codes, both in terms of speed and error correction capabilities, has become a hot topic within the QEC literature.
@@
Surface codes are one of the most promising families of codes for constructing primitive fault-tolerant quantum computers in the near-term [25, 31, 45, 46, 47, 48, 49]. This family of topological codes presents the benefit of being implementable in two-dimensional grids of qubits with local syndrome measurements, or check operators in the surface code terminology, while presenting a high tolerance to quantum noise [46]. Considering that many of the physical platforms being considered for constructing quantum processors such as quantum dots, cold atoms or even the mainstream superconducting qubits have architectural restrictions limiting qubit connectivity; surface codes represent a perfect candidate for implementing QEC using those technologies. Recently, major breakthroughs in the field of quantum error correction have been achieved with the first successful experimental implementations of surface codes over superconducting qubit and neutral atom processors [50, 51, 52]. The result by Google Quantum AI is specially relevant due to the fact that it experimentally proves that QEC improves its performance when the distance of the surface code increases [51]. Importantly, the recent experiments by Harvard and QuEra also showed such scaling in the application of a transversal CNOT operation as well as other important results for QEC [52].
@@
In this sense, designing decoders for surface codes is a really important task for near-term quantum computers. At the time of writing, there are many potential methods in order to perform the inference of the error from the syndrome data, each of them with their own strong and weak points. Generally, the performance versus decoding complexity trade-off stands for those methods and, therefore, each of them is a potential candidate for being the one selected for experimental quantum error correction depending on the specific needs of the system to be error corrected. Due to the fact that surface code decoding is a relevant and timely topic, the main aim of this tutorial is twofold. The first is to provide a compilation and a comprehensive description of the main decoders for surface codes, while the second is to offer a comparison between those methods in terms of decoding complexity and performance. With such scope in mind, we first provide an introductory section describing the basic notions of stabilizer code theory in Section 2 so that surface codes can be introduced in Section 3. We follow by discussing the noise sources in quantum computers and the way they are modelled in Section 4. These sections serve as preliminary background in order to understand the decoding methods that are discussed in the core of the review, Section 5. We describe the most popular surface code decoding algorithms: the Minimum-Weight Perfect Matching (MWPM) decoder, the Union-Find (UF) decoder, the Belief Propagation (BP) decoder and the Tensor Network (TN) decoder. We present their functioning in a comprehensive manner and later discuss not only their performance under depolarizing and biased noise via simulations but also their computational complexity. In addition, we also discuss other decoding methods proposed through the literature for general surface (topological) codes. Also, we review existing and publicly available software implementations of the discussed decoding methods. We then provide a discussion section, Section 6, where we compare the decoders for rotated planar codes and provide an overview of the challenges in decoding surface codes.
@@ 2 Background
In stark contrast, the building block of quantum computers are the elements referred as qubits. Those quantum mechanical systems, are two-level quantum systems that allow coherent superpositions of them. In this sense, a qubit can be described as a vector in a two-dimensional complex Hilbert space, $\mathcal{H}_{2}$ [15]:
@@
This and other properties of quantum mechanics such as entanglement allow for a series of advantages in computation (speedups in algorithm complexity [3, 4]), or in communications (supperadditivity of the quantum channel capacity [53]), among others. Nevertheless, such promises are put in question by the inherent noise present in these quantum mechanical systems. As good as it is to make use of quantum unique properties such as superposition or entanglement, quantum noise does also follow different laws and, thus, it is somewhat different to the noise in classical computers/communications. Classical bit noise can be summarized to flips, a logical operation which would turn the 0 into $1$ and vice versa. We refer to Section 4 for understanding quantum noise in general, but as it will be seen there, for the case of a single qubit the noise can be described by means of Pauli channels. The Pauli channel refers to a stochastic model where a qubit can suffer from bit-flips, phase-flips or bit-and-phase-flips each with some probability. With this consideration in mind, these noise effects are described by the elements of the Pauli group [15], $\mathcal{G}_{1}$, whose generators are the Pauli matrices $\langle X,Y,Z\rangle$. Those operators transform the state of an arbitrary qubit as in eq.2 in the following way:
@@
Note that the the $Y$ operator not only does perform a bit-and-phase-flip operation on the arbitrary quantum state but also changes its overall phase. However, neglecting the global phase has no observable physical consequences and, thus, it is often completely ignored from the point of view of quantum error correction [54].
Dealing with these noise processes is a task of vital importance if complex quantum algorithms are meant to be executed reliably. In this context, there are two main approaches to deal with noisy quantum computations: quantum error mitigation (QEM) and quantum error correction (QEC). The two approaches have shown to be complementary with recent papers proposing schemes such as distance scaled zero-noise extrapolation (DSZNE) [55]. QEM attempts to evaluate accurate expectation values of physical observables of interest by using noisy qubits and quantum circuits [55, 56], while the main objective of QEC is to obtain qubits and computations that experience error rates that are arbitrarily low. For example, it is considered that a logical error probability in the order of $10^{-14}-10^{-15}$ is required to run Shor’s algorithm fault-tolerantly [25]. There are many approaches to quantum error correction, but the general idea behind QECCs is to protect the information of a number of qubits $k$, named logical qubits, within a larger number of qubits $n$, named physical qubits in a way that makes the whole system tolerant to a number of errors. Within these QECCs, many lay within the framework named Quantum Stabilizer Coding [30]. Stabilizer codes allow for a direct mapping of many classical methods into the quantum spectrum, making them very useful [30]. Since the surface code belongs to the family of quantum stabilizer codes, it is pertinent to cover the basics of such QEC constructions.
@@ 2.1 Stabilizer Codes
if $E\in\mathcal{S}\subset\mathcal{G}_{n}$. In this sense, there will be different error operators that share the same error syndrome that affects the encoded quantum state in a similar manner. This phenomenon is usually termed as degeneracy. The concept of error degeneracy has the consequence that the Pauli space that represents all possible error operators is not just partitioned into syndrome cosets, but also into degenerate error cosets [^8] Note that degeneracy is somehow different in the entanglement-assisted paradigm [57] [41]. Specifically, the Pauli group is partitioned in $2^{n-k}$ cosets that share error syndrome, and each of those cosets will be partitioned in to $2^{2k}$ cosets that contain $2^{n-k}$ errors that are degenerate among them [53, 41]. How degenerate a quantum code is depends on the difference between the weight of its stabilizer generators and its distance. If $w(M_{k})<<d$, $\forall k\in{0,\ldots,n-k}$, where $M_{k}\in\mathcal{S}$ denotes a stabilizer generator and $w$ denotes the weight, then each logical coset (equivalence class) will contain many operators of the same weight and the code will be highly degenerate. In cases where $w(M_{k})=d$, the code will be non-degenerate.
In summary, the checks give us a partial information of the error operator that corrupted the encoded information. Since the aim of quantum error correction is to recover the noiseless quantum state, an estimate of the channel error must be obtained so that the noisy state can be corrected. The process of estimating the quantum error from the measured syndrome is named decoding. Once a guess of the error, $\hat{E}\in\mathcal{G}_{n}$, is obtained by the decoder, it will be operated on the code. If the estimation turns out to be correct, the noisy quantum state will be successfully corrected since the elements of the Pauli group are self-inverse. Moreover, if the estimated error is not the exact element of the Pauli group but it belongs to the same degenerate coset, the correction will also be successful [41]. Finally, whenever the estimated error does not fulfill any of those two cases, the correction operation will result in a non-trivial action on the logical qubits encoded in the state, implying that the error correction method has failed.